A vessel at 1000 K contains CO2 with a pressure of 0.5 atm. Some of the CO2 is converted into CO on the addition of graphite. If the total pressure at equilibrium is 0.8 atm, the value of K is ?
A 0.18 atm
B 1.8 atm
C 3 atm
D 0.3 atm

Grade:11

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

The answer is option b

Regards
Vikas TU
3 years ago
initially        0.5    0             0
at eq'm     0.5-x    0           2x
As solids do not contribute to pressure.
Given total pressure= 0.5-x+2x=0.8
x=0.3
Thus 
pco2 = 0.5-0.3 = 0.2 
pco = 2*0.3 
So eq'm constant will be:
kp = p(co)^2 / (pco2) 
= 1.8